# A Benchmark for Multi-modal Foundation Models on Low-level Vision: from Single Images to Pairs

### v1

@@ Abstract
The rapid development of Multi-modality Large Language Models (MLLMs) has navigated a paradigm shift in computer vision, moving towards versatile foundational models. However, evaluating MLLMs in low-level visual perception and understanding remains a yet-to-explore domain. To this end, we design benchmark settings to emulate human language responses related to low-level vision: the low-level visual perception (A1) via visual question answering related to low-level attributes (e.g. clarity, lighting); and the low-level visual description (A2), on evaluating MLLMs for low-level text descriptions. Furthermore, given that pairwise comparison can better avoid ambiguity of responses and has been adopted by many human experiments, we further extend the low-level perception-related question-answering and description evaluations of MLLMs from single images to image pairs. Specifically, for perception (A1), we carry out the LLVisionQA dataset, comprising 2,990 single images and 1,999 image pairs each accompanied by an open-ended question about its low-level features; for description (A2), we propose the LLDescribe dataset, evaluating MLLMs for low-level descriptions on 499 single images and 450 pairs. Additionally, we evaluate MLLMs on assessment (A3) ability, i.e. predicting score, by employing a softmax-based approach to enable all MLLMs to generate quantifiable quality ratings, tested against human opinions in 7 image quality assessment (IQA) datasets. With 24 MLLMs under evaluation, we demonstrate that several MLLMs have decent low-level visual competencies on single images, but only GPT-4V exhibits higher accuracy on pairwise comparisons than single image evaluations (like humans). We hope that our benchmark will motivate further research into uncovering and enhancing these nascent capabilities of MLLMs. Datasets will be available at https://github.com/Q-Future/Q-Bench.
@@
The emergent large language models (LLMs) such as ChatGPT and Bard, as well as their excellent open-source counterparts (e.g., LLaMA [1], MPT [2]), have served as powerful general-purpose assistants, which opens a new era for artificial intelligence (AI) from targeting specific tasks towards general intelligence. Following the advancements of LLMs, multi-modality large language models (MLLMs), as represented by LLaVA [3], MiniGPT-4 [4], InstructBLIP [5], and Otter [6], have brought exciting progresses on the vision field as well. They are capable of providing robust general-level abilities on visual perception/understanding and can even seamlessly dialog and interact with humans through natural language. While such abilities of MLLMs have been validated on several vision-language tasks such as image captioning [7], visual question answering [8], cross-modality grounding [9], and traditional vision tasks such as image classification or segmentation [10], most attention is paid to the high-level perception and understanding of visual contents. Meanwhile, the ability of MLLMs remains unclear on low-level visual perception and understanding, which play significant roles in image quality assessment (IQA) [11, 12] and its associated tasks on perceiving visual distortions (noises, blurs) [13, 14], and other low-level attributes (color, lighting, composition, style, etc) [15] that may relate to aesthetics of natural photos [16] as well as human preferences on emerging computer-graphics generated [17] or AI-generated images [18, 19]. These low-level visual abilities are strongly associated with a wide range of applications, such as recommendation [20], guidance on camera systems [21], or visual quality enhancement [22]. Henceforth, it is crucial to evaluate these general-purpose foundation models in low-level visual perception and understanding, to relieve extensive human resources on giving feedback to every specific low-level task.
In this paper, we propose the first systematic benchmark Q-Bench to measure the low-level visual abilities of MLLMs, which is constructed around a key question:
@@
Although the above two capabilities essentially emulate human perception of low-level vision, they still miss some key capabilities of humans. For example, regarding Fig. 1 (a), some people may consider its clarity to be average, while others may deem it poor, while neither opinion should be considered incorrect; instead, everyone would agree that Fig. 1 (b) is clearer than Fig. 1 (a). On the other hand, for those who regard both Fig. 1 (b) and Fig. 1 (a) as blurry, comparing the clarity between the pair can also provide additional valuable information. Noticing these issues, lots of recent subjective studies [22, 23, 24] have adopted the a juxtaposition-based paradigm, that is, collecting human opinions by comparing a pair of images. Based on these insights and recent progresses on MLLMs [25, 26, 27, 28] that officially support more than one images as inputs, we further explore whether MLLMs can similarly emulate respective human capabilities:
@@
I-1 LLVisionQA Benchmark Dataset
To evaluate the low-level perception ability (A1) on various low-level attributes under diverse circumstances, we construct the LLVisionQA dataset, including 2,990 single images and 1,999 image pairs from 10 diverse sources. Aligned with existing practices [29, 30], each single image or image pair in LLVisionQA is equipped with a question, alongside a correct answer and false candidate answers. Specifically, we design three diverse types of questions: Yes-or-No questions, What questions, and How questions. Moreover, we divide low-level concerns for single images into four quadrants, via two axes: (1) distortions (blur, noises, etc) vs other low-level attributes (color, lighting, composition, etc) [31]. (2) global perception (e.g., sharpness of the whole picture) vs local content-related in-context perception (e.g., whether the red flower is in focus) [32]. On the other hand, we separate the low-level concerns for image pairs into four sub-categories: (1) distortions vs other low-level attributes (similar as above). (2) comparison (e.g., which image is clearer) vs joint analysis (e.g., are both images underexposure). With three types of questions and divided concerns, the proposed LLVisionQA dataset provides a holistic benchmark for the low-level perception abilities of MLLMs on both single images and pairs.
I-2 LLDescribe Benchmark Dataset
For the description ability (A2), given that the output description is expected to be open-ended, we propose the LLDescribe dataset by inviting experts with rich experience in the low-level vision field to write long golden low-level descriptions (average 58 words per description) for 499 single images and 450 image pairs. The long golden low-level descriptions then serve as the reference texts for the single-modal GPT to evaluate MLLM output descriptions. To ensure the evaluation is comprehensive, the quality of MLLM descriptions is evaluated through three dimensions: completeness (punish missing information), preciseness (punish outputs controversial with reference), as well as relevance (punish outputs irrelevant to low-level attributes). With golden descriptions and the multi-dimensional evaluation process participated by GPT, we comprehensively evaluate the low-level description ability of MLLMs.
@@ I-3 IQA Benchmark
For the assessment ability, we utilize plenty of existing IQA databases [11, 33, 18, 17, 12, 34] that focus on various low-level appearances of images, to benchmark MLLMs within conventional IQA settings. Specifically, we notice that MLLMs encounter difficulties in providing sufficiently quantifiable outputs, whether instructed to directly rate with texts or provide numerical outputs. To solve this challenge, we propose to extract the softmax pooling result on the logits of the two most frequent tokens (good and poor) under the response template of MLLMs (Fig. 2 (A3)) as their quality predictions. Our studies prove that the proposed softmax-based strategy is generally better correlated with human perception than direct token outputs of MLLMs (via argmax), which bridges between these emergent MLLMs and the traditional IQA task settings. Under this strategy, we evaluate all MLLMs on their precise assessment ability by measuring the correlations between their predictions and human opinion scores in various IQA databases. Furthermore, we propose a prompt-ensemble approach to help boost the IQA performance of MLLMs with the softmax-based strategy.
This work is a substantially extended version of our earlier conference publication [35]. Compared with the conference version, we bring three major changes: (1) Most importantly, we extend the perception and description tasks from single images to image pairs, which provides a more comprehensive benchmark for MLLMs on emulating human low-level visual understanding ability. (2) We update the benchmark with the latest popular MLLMs (evaluated MLLMs increased from 15 to 24), providing a review of the development for MLLMs on low-level vision. (3) We further propose a simple yet effective prompt-ensemble approach, which can help boost the zero-shot performance of MLLMs on the assessment task.
In summary, we systematically explore the potential of MLLMs on three low-level visual abilities: perception, description, and assessment. The three realms compose into the proposed Q-Bench, a MLLM benchmark on low-level visual tasks. Our contributions can be summarized as three-fold:
We build a benchmark for MLLMs on low-level perception ability. To achieve this, we construct a first-of-its-kind balanced and comprehensive LLVisionQA dataset with 2,990 single images and 1,999 image pairs with one low-level-related question-answer pair for each image. The LLVisionQA dataset includes three question types and multiple low-level concerns to ensure diversity.
We define a benchmark process to evaluate the low-level description ability of MLLMs, including an LLDescription dataset of 499 single images and 450 image pairs with expert-labeled long golden quality descriptions, and a GPT-assisted evaluation to rate MLLM-descriptions in terms of completeness, preciseness, and relevance compared with golden descriptions.
@@
| Type | Source Dataset | LLVisionQA | LLDescribe | Full Dataset Size |
@@
| In-the-wild | KONiQ-10K [11] | 600 | 200 | 10,073 |
| SPAQ [12] | 800 | 200 | 11,125 | |
| LIVE-FB [37] | 300 | 50 | 39,810 | |
| LIVE-itw [38] | 300 | 50 | 1,169 | |
@@
| AGIQA-3K [18] | 198 | 80 | 2,982 | |
| ImageRewardDB [19] | 194 | 29 | excluded in (A3) | |
@@
| LIVEMultiDistortion [39] | 15 | 10 | excluded in (A3) | |
| Corrupted COCO [7] | 302 | 50 | excluded in (A3) | |
| Corresponding Task in Q-Bench | (A1) Perception | (A2) Description | (A3) Assessment | |
@@ II-A1 Focusing on Low-level Visual Abilities of MLLMs
Unlike existing MLLM benchmarks [40, 29, 30] that aim at all-round abilities, the tasks in Q-Bench are constrained with two basic principles: a) Requiring perception and/or understanding on low-level attributes of images; b) Not requiring reasoning (i.e. why) or outside knowledge [41]. We adhere to the principles in designing the perception, description, and assessment tasks, making the proposed Q-Bench a focused reflection on the low-level visual abilities of MLLMs.
@@ II-A2 Covering Diverse Low-level Appearances
To cover diverse low-level appearances, we collect multi-sourced images for each task, as depicted in Table I. Among all images in the perception and description tasks, two-thirds are in-the-wild images directly collected from social media posts, smartphones, or professional photography. The rest one-third images are collected after various artificial distortions, or via generative processes (CGI, AIGC). Furthermore, we employ k-means clustering for the low-level attribute indicators to certify that the sub-sampled images retain high diversity. In the assessment task, full images of 7 IQA datasets within all three source types are evaluated through traditional IQA metrics. The diverse and multiple sources of images morph the Q-Bench into a holistic and balanced benchmark to fairly evaluate low-level-related abilities.
@@ II-A3 Extending from Single Images to Image Pairs
Evaluating image pairs allows for direct comparison and joint analysis of low-level attributes, which can highlight subtle differences or similarities that might not be evident when images are viewed in isolation. Humans are good at comparing, therefore we believe it is also important to benchmark the low-level visual perception and understanding ability of MLLMs on image pairs. Thus we extend the benchmark (only including single images) in our conference version [35] with image pairs to simulate more complex visual tasks that mirror real-world scenarios and challenge the MLLMs to process and compare multiple visual inputs simultaneously.
@@ II-B Benchmark on Low-level Perception Ability
In the first task of Q-Bench, we evaluate the low-level perception ability of MLLMs to examine whether they can answer simple natural queries related to low-level attributes. For this purpose, we first collect 2,990 single images (I) from multiple sources (see Table I) with diverse low-level concerns, from which we collect 1,999 image pairs (I’) as well. All image pairs are different from each other but may have one repeated image across different pairs. Then, we collect one low-level-related question (Q), one correct answer to the question (C), and 1-3 candidate false answers (F) for each single image or image pair. The 2,990 (I,Q,C,F) and 1,999 (I’,Q,C,F) tuples compose into the LLVisionQA dataset (as illustrated in Fig. 3), the first visual question answering (VQA) dataset in the low-level computer vision field. Specifically, the questions in LLVisionQA cover four quadrants of distinct low-level concerns and three question types. After constructing the dataset, the (I,Q,C,F) are together fed into MLLMs for evaluation, while their outputs are further examined by GPT to judge correctness. The details are elaborated as follows.
@@
Axis 1: Distortions vs Other Low-level Attributes. The primary axis differentiates two categories of low-level perceptual attributes: 1) technical distortions [13], seen as the low-level characteristics that directly degrade the quality of images [37], and 2) aesthetic-related other low-level attributes [15, 42] which are discernible to human perception and evoke varied emotions. Several studies [43, 37, 31] follow this paradigm and categorize them through a relative golden standard, that whether the attributes directly improve or degrade picture quality (YesDistortions; NoOthers).
Axis 2: Global Perception vs Local In-context Perception. In recent research on low-level vision, it is observed that human perceptions of low-level visuals often intertwine with higher-level contextual comprehension [32, 44, 45, 46]. For instance, a clear sky might lack complex textures yet display exceptional clarity. Furthermore, localized low-level appearances can deviate from their overall counterparts, as observed by [47, 48]. Acknowledging these differences, we curate local in-context perception (Fig. 3 right top) questions, that require MLLMs to grasp the content or other context to answer correctly, while other questions are categorized as global perception (Fig. 3 left top).
@@
Axis 2: Compare vs Joint. This dual approach mimics human visual perception more closely. Humans often use both comparison (looking at differences and similarities) and joint analysis (perceiving images in a unified context) when viewing images. The comparison highlights the differences and similarities between the two images, which is the key component of the full-reference IQA tasks [49] and other low-level enhancement evaluation tasks [50, 51]. The joint analysis, on the other hand, looks at the images as a combined entity to understand the overall context or to detect patterns that emerge only when the images are considered together.
@@
In the LLVisionQA dataset, we curate three question types, Yes-or-No, What, and How to simulate multiple query forms from humans. The details of the three question types are defined as follows.
@@
Type 2: What Questions. Despite Yes-or-No judgments, the what questions are also a common type of queries in recent MLLM benchmarks such as [30]. In Q-bench, they classify low-level attributes in pictures (e.g., What distortion occurs in the image?), or associated context given specific low-level appearances (for in-context perception questions, e.g., Which object in the image is under-exposed?). Unlike Yes-or-No questions, the What questions examine more comprehensive low-level attribute understanding of MLLMs, by requiring correct perception on multiple attributes.
Type 3: How Questions. Despite the two common types, we also include a special type, the How questions, to cover non-extreme appearances [14] of low-level attribute dimensions into our benchmark, as an extension to Yes-or-No questions. As shown in Fig. 3, we can query How is the clarity of the image? for the image with both clear and blurry areas, and answer with Medium. With this special question type, we broaden the Q-bench into finer-grained low-level perception.
@@
The correct answer has been shuffled and finally uniformly distributed among all choices (A/B/C/D). Moreover, while traditional visual question answering [8, 41] tasks typically employ traditional language metrics (BLEU-4, CIDEr) to compare performance, as observed by recent studies [52] and validated by us, most MLLMs cannot consistently provide outputs on instructed formats. Given the question above, different MLLMs may reply “A.”, “High”, “The clarity of the image is high.”, “The image is of high clarity.” (all correct), which are difficult to be exhaustively-included under traditional metrics. To solve this problem, we design, validate, and employ a 5-round GPT-assisted evaluation process inspired by [29]. Under this process, the question, correct answers, and MLLM replies are fed into GPT for evaluation.

### v2

@@ Abstract
The rapid development of Multi-modality Large Language Models (MLLMs) has navigated a paradigm shift in computer vision, moving towards versatile foundational models. However, evaluating MLLMs in low-level visual perception and understanding remains a yet-to-explore domain. To this end, we design benchmark settings to emulate human language responses related to low-level vision: the low-level visual perception (A1) via visual question answering related to low-level attributes (e.g. clarity, lighting); and the low-level visual description (A2), on evaluating MLLMs for low-level text descriptions. Furthermore, given that pairwise comparison can better avoid ambiguity of responses and has been adopted by many human experiments, we further extend the low-level perception-related question-answering and description evaluations of MLLMs from single images to image pairs. Specifically, for perception (A1), we carry out the LLVisionQA+ dataset, comprising 2,990 single images and 1,999 image pairs each accompanied by an open-ended question about its low-level features; for description (A2), we propose the LLDescribe+ dataset, evaluating MLLMs for low-level descriptions on 499 single images and 450 pairs. Additionally, we evaluate MLLMs on assessment (A3) ability, i.e. predicting score, by employing a softmax-based approach to enable all MLLMs to generate quantifiable quality ratings, tested against human opinions in 7 image quality assessment (IQA) datasets. With 24 MLLMs under evaluation, we demonstrate that several MLLMs have decent low-level visual competencies on single images, but only GPT-4V exhibits higher accuracy on pairwise comparisons than single image evaluations (like humans). We hope that our benchmark will motivate further research into uncovering and enhancing these nascent capabilities of MLLMs. Datasets will be available at https://github.com/Q-Future/Q-Bench.
@@
The emergent large language models (LLMs) such as ChatGPT and Bard, as well as their excellent open-source counterparts (e.g., LLaMA [llama], MPT [mpt]), have served as powerful general-purpose assistants, which opens a new era for artificial intelligence (AI) from targeting specific tasks towards general intelligence. Following the advancements of LLMs, multi-modality large language models (MLLMs), as represented by LLaVA [llava], MiniGPT-4 [minigpt4], InstructBLIP [iblip], and Otter [otter], have brought exciting progresses on the vision field as well. They are capable of providing robust general-level abilities on visual perception/understanding and can even seamlessly dialog and interact with humans through natural language. While such abilities of MLLMs have been validated on several vision-language tasks such as image captioning [cococaps], visual question answering [cocovqa], cross-modality grounding [kosmos2], and traditional vision tasks such as image classification or segmentation [lai2023lisa], most attention is paid to the high-level perception and understanding of visual contents. Meanwhile, the ability of MLLMs remains unclear on low-level visual perception and understanding, which play significant roles in image quality assessment (IQA) [koniq, spaq] and its associated tasks on perceiving visual distortions (noises, blurs) [koniqplusplus, wu2023explainable], and other low-level attributes (color, lighting, composition, style, etc) [aadb] that may relate to aesthetics of natural photos [avaiaa] as well as human preferences on emerging computer-graphics generated [zhang2023subjective] or AI-generated images [agiqa3k, imagereward]. These low-level visual abilities are strongly associated with a wide range of applications, such as recommendation [wu2023dover], guidance on camera systems [irpotential], or visual quality enhancement [lpips]. Henceforth, it is crucial to evaluate these general-purpose foundation models in low-level visual perception and understanding, to relieve extensive human resources on giving feedback to every specific low-level task.
In this paper, we propose the first systematic benchmark Q-Bench+ to measure the low-level visual abilities of MLLMs, which is constructed around a key question:
@@
Although the above two capabilities essentially emulate human perception of low-level vision, they still miss some key capabilities of humans. For example, regarding Fig. 1 (a), some people may consider its clarity to be average, while others may deem it poor, while neither opinion should be considered incorrect; instead, everyone would agree that Fig. 1 (b) is clearer than Fig. 1 (a). On the other hand, for those who regard both Fig. 1 (b) and Fig. 1 (a) as blurry, comparing the clarity between the pair can also provide additional valuable information. Noticing these issues, lots of recent subjective studies [lpips, pieapp, pipal] have adopted the a juxtaposition-based paradigm, that is, collecting human opinions by comparing a pair of images. Based on these insights and recent progresses on MLLMs [emu2, bakllava, geminipro, openai2023gpt4] that officially support more than one images as inputs, we further explore whether MLLMs can similarly emulate respective human capabilities:
@@
I-1 LLVisionQA+ Benchmark Dataset
To evaluate the low-level perception ability (A1) on various low-level attributes under diverse circumstances, we construct the LLVisionQA+ dataset, including 2,990 single images and 1,999 image pairs from 10 diverse sources. Aligned with existing practices [mmbench, emabench], each single image or image pair in LLVisionQA+ is equipped with a question, alongside a correct answer and false candidate answers. Specifically, we design three diverse types of questions: Yes-or-No questions, What questions, and How questions. Moreover, we divide low-level concerns for single images into four quadrants, via two axes: (1) distortions (blur, noises, etc) vs other low-level attributes (color, lighting, composition, etc) [atqa]. (2) global perception (e.g., sharpness of the whole picture) vs local content-related in-context perception (e.g., whether the red flower is in focus) [sfa]. On the other hand, we separate the low-level concerns for image pairs into four sub-categories: (1) distortions vs other low-level attributes (similar as above). (2) comparison (e.g., which image is clearer) vs joint analysis (e.g., are both images underexposure). With three types of questions and divided concerns, the proposed LLVisionQA+ dataset provides a holistic benchmark for the low-level perception abilities of MLLMs on both single images and pairs.
I-2 LLDescribe+ Benchmark Dataset
For the description ability (A2), given that the output description is expected to be open-ended, we propose the LLDescribe+ dataset by inviting experts with rich experience in the low-level vision field to write long golden low-level descriptions (average 58 words per description) for 499 single images and 450 image pairs. The long golden low-level descriptions then serve as the reference texts for the single-modal GPT to evaluate MLLM output descriptions. To ensure the evaluation is comprehensive, the quality of MLLM descriptions is evaluated through three dimensions: completeness (punish missing information), preciseness (punish outputs controversial with reference), as well as relevance (punish outputs irrelevant to low-level attributes). With golden descriptions and the multi-dimensional evaluation process participated by GPT, we comprehensively evaluate the low-level description ability of MLLMs.
@@ I-3 IQA Benchmark
For the assessment ability, we utilize plenty of existing IQA databases [koniq, kadid, agiqa3k, zhang2023subjective, spaq, livechallenge] that focus on various low-level appearances of images, to benchmark MLLMs within conventional IQA settings. Specifically, we notice that MLLMs encounter difficulties in providing sufficiently quantifiable outputs, whether instructed to directly rate with texts or provide numerical outputs. To solve this challenge, we propose to extract the softmax pooling result on the logits of the two most frequent tokens (good and poor) under the response template of MLLMs (Fig. 2 (A3)) as their quality predictions. Our studies prove that the proposed softmax-based strategy is generally better correlated with human perception than direct token outputs of MLLMs (via argmax), which bridges between these emergent MLLMs and the traditional IQA task settings. Under this strategy, we evaluate all MLLMs on their precise assessment ability by measuring the correlations between their predictions and human opinion scores in various IQA databases. Furthermore, we propose a prompt-ensemble approach to help boost the IQA performance of MLLMs with the softmax-based strategy.
This work is a substantially extended version of our earlier conference publication [wu2024qbench]. Compared with the conference version, we bring three major changes: (1) Most importantly, we extend the perception and description tasks from single images to image pairs, which provides a more comprehensive benchmark for MLLMs on emulating human low-level visual understanding ability. (2) We update the benchmark with the latest popular MLLMs (evaluated MLLMs increased from 15 to 24), providing a review of the development for MLLMs on low-level vision. (3) We further propose a simple yet effective prompt-ensemble approach, which can help boost the zero-shot performance of MLLMs on the assessment task.
In summary, we systematically explore the potential of MLLMs on three low-level visual abilities: perception, description, and assessment. The three realms compose into the proposed Q-Bench+, a MLLM benchmark on low-level visual tasks. Our contributions can be summarized as three-fold:
We build a benchmark for MLLMs on low-level perception ability. To achieve this, we construct a first-of-its-kind balanced and comprehensive LLVisionQA+ dataset with 2,990 single images and 1,999 image pairs with one low-level-related question-answer pair for each image. The LLVisionQA+ dataset includes three question types and multiple low-level concerns to ensure diversity.
We define a benchmark process to evaluate the low-level description ability of MLLMs, including an LLDescription+ dataset of 499 single images and 450 image pairs with expert-labeled long golden quality descriptions, and a GPT-assisted evaluation to rate MLLM-descriptions in terms of completeness, preciseness, and relevance compared with golden descriptions.
@@
| Type | Source Dataset | LLVisionQA+ | LLDescribe+ | Full Dataset Size |
@@
| In-the-wild | KONiQ-10K [koniq] | 600 | 200 | 10,073 |
| SPAQ [spaq] | 800 | 200 | 11,125 | |
| LIVE-FB [paq2piq] | 300 | 50 | 39,810 | |
| LIVE-itw [clive] | 300 | 50 | 1,169 | |
@@
| AGIQA-3K [agiqa3k] | 198 | 80 | 2,982 | |
| ImageRewardDB [imagereward] | 194 | 29 | excluded in (A3) | |
@@
| LIVEMultiDistortion [livemultipledistortions] | 15 | 10 | excluded in (A3) | |
| Corrupted COCO [cococaps] | 302 | 50 | excluded in (A3) | |
| Corresponding Task in Q-Bench+ | (A1) Perception | (A2) Description | (A3) Assessment | |
@@ II-A1 Focusing on Low-level Visual Abilities of MLLMs
Unlike existing MLLM benchmarks [seedbench, mmbench, emabench] that aim at all-round abilities, the tasks in Q-Bench+ are constrained with two basic principles: a) Requiring perception and/or understanding on low-level attributes of images; b) Not requiring reasoning (i.e. why) or outside knowledge [okvqa]. We adhere to the principles in designing the perception, description, and assessment tasks, making the proposed Q-Bench+ a focused reflection on the low-level visual abilities of MLLMs.
@@ II-A2 Covering Diverse Low-level Appearances
To cover diverse low-level appearances, we collect multi-sourced images for each task, as depicted in Table I. Among all images in the perception and description tasks, two-thirds are in-the-wild images directly collected from social media posts, smartphones, or professional photography. The rest one-third images are collected after various artificial distortions, or via generative processes (CGI, AIGC). Furthermore, we employ k-means clustering for the low-level attribute indicators to certify that the sub-sampled images retain high diversity. In the assessment task, full images of 7 IQA datasets within all three source types are evaluated through traditional IQA metrics. The diverse and multiple sources of images morph the Q-Bench+ into a holistic and balanced benchmark to fairly evaluate low-level-related abilities.
@@ II-A3 Extending from Single Images to Image Pairs
Evaluating image pairs allows for direct comparison and joint analysis of low-level attributes, which can highlight subtle differences or similarities that might not be evident when images are viewed in isolation. Humans are good at comparing, therefore we believe it is also important to benchmark the low-level visual perception and understanding ability of MLLMs on image pairs. Thus we extend the benchmark (only including single images) in our conference version [wu2024qbench] with image pairs to simulate more complex visual tasks that mirror real-world scenarios and challenge the MLLMs to process and compare multiple visual inputs simultaneously.
@@ II-B Benchmark on Low-level Perception Ability
In the first task of Q-Bench+, we evaluate the low-level perception ability of MLLMs to examine whether they can answer simple natural queries related to low-level attributes. For this purpose, we first collect 2,990 single images (I) from multiple sources (see Table I) with diverse low-level concerns, from which we collect 1,999 image pairs (I’) as well. All image pairs are different from each other but may have one repeated image across different pairs. Then, we collect one low-level-related question (Q), one correct answer to the question (C), and 1-3 candidate false answers (F) for each single image or image pair. The 2,990 (I,Q,C,F) and 1,999 (I’,Q,C,F) tuples compose into the LLVisionQA+ dataset (as illustrated in Fig. 3), the first visual question answering (VQA) dataset in the low-level computer vision field. Specifically, the questions in LLVisionQA+ cover four quadrants of distinct low-level concerns and three question types. After constructing the dataset, the (I,Q,C,F) are together fed into MLLMs for evaluation, while their outputs are further examined by GPT to judge correctness. The details are elaborated as follows.
@@
Axis 1: Distortions vs Other Low-level Attributes. The primary axis differentiates two categories of low-level perceptual attributes: 1) technical distortions [koniqplusplus], seen as the low-level characteristics that directly degrade the quality of images [paq2piq], and 2) aesthetic-related other low-level attributes [aadb, clipiaa] which are discernible to human perception and evoke varied emotions. Several studies [nima, paq2piq, atqa] follow this paradigm and categorize them through a relative golden standard, that whether the attributes directly improve or degrade picture quality (YesDistortions; NoOthers).
Axis 2: Global Perception vs Local In-context Perception. In recent research on low-level vision, it is observed that human perceptions of low-level visuals often intertwine with higher-level contextual comprehension [sfa, rfugc, wu2022fastervqa, qalign]. For instance, a clear sky might lack complex textures yet display exceptional clarity. Furthermore, localized low-level appearances can deviate from their overall counterparts, as observed by [fastvqa, pvq]. Acknowledging these differences, we curate local in-context perception (Fig. 3 right top) questions, that require MLLMs to grasp the content or other context to answer correctly, while other questions are categorized as global perception (Fig. 3 left top).
@@
Axis 2: Compare vs Joint. This dual approach mimics human visual perception more closely. Humans often use both comparison (looking at differences and similarities) and joint analysis (perceiving images in a unified context) when viewing images. The comparison highlights the differences and similarities between the two images, which is the key component of the full-reference IQA tasks [hore2010image] and other low-level enhancement evaluation tasks [zhang2021no, zhang2022no]. The joint analysis, on the other hand, looks at the images as a combined entity to understand the overall context or to detect patterns that emerge only when the images are considered together.
@@
In the LLVisionQA+ dataset, we curate three question types, Yes-or-No, What, and How to simulate multiple query forms from humans. The details of the three question types are defined as follows.
@@
Type 2: What Questions. Despite Yes-or-No judgments, the what questions are also a common type of queries in recent MLLM benchmarks such as [emabench]. In Q-bench+, they classify low-level attributes in pictures (e.g., What distortion occurs in the image?), or associated context given specific low-level appearances (for in-context perception questions, e.g., Which object in the image is under-exposed?). Unlike Yes-or-No questions, the What questions examine more comprehensive low-level attribute understanding of MLLMs, by requiring correct perception on multiple attributes.
Type 3: How Questions. Despite the two common types, we also include a special type, the How questions, to cover non-extreme appearances [wu2023explainable] of low-level attribute dimensions into our benchmark, as an extension to Yes-or-No questions. As shown in Fig. 3, we can query How is the clarity of the image? for the image with both clear and blurry areas, and answer with Medium. With this special question type, we broaden the Q-bench+ into finer-grained low-level perception.
@@
The correct answer has been shuffled and finally uniformly distributed among all choices (A/B/C/D). Moreover, while traditional visual question answering [cocovqa, okvqa] tasks typically employ traditional language metrics (BLEU-4, CIDEr) to compare performance, as observed by recent studies [mplugowl] and validated by us, most MLLMs cannot consistently provide outputs on instructed formats. Given the question above, different MLLMs may reply “A.”, “High”, “The clarity of the image is high.”, “The image is of high clarity.” (all correct), which are difficult to be exhaustively-included under traditional metrics. To solve this problem, we design, validate, and employ a 5-round GPT-assisted evaluation process inspired by [mmbench]. Under this process, the question, correct answers, and MLLM replies are fed into GPT for evaluation.